OpenStudy (anonymous):

Simplifysquare root of 2 multiplied by the cube root of 2. 2 to the power of 1 over 6 2 to the power of 2 over 3 2 to the power of 5 over 6 2 to the power of 7 over 6

OpenStudy (bibby):

recall that roots can be written as fractional exponents like so: \(\huge \sqrt{x}=x^{\frac{1}{2}}\) \(\huge \sqrt[4]{x}=x^{\frac{1}{4}}\)

OpenStudy (bibby):

also recall that \(x^m*x^n=x^{m+n}\)

OpenStudy (anonymous):

So whats the answer though? i don't understand.

OpenStudy (wolf1728):

X^(1/2) * x^(1/3) = x ^ (1/2 + 1/3) Does that help?

OpenStudy (anonymous):

Is it B?

OpenStudy (wolf1728):

No- what is 1/2 plus 1/3 ?

OpenStudy (anonymous):

5/6 so C.

OpenStudy (wolf1728):

Very good !!!!!!

OpenStudy (anonymous):

Sorry.. I'm not too good with math, and my lack of sleep isnt really helping much either. But thank you so much for your help!

OpenStudy (wolf1728):

glad to help out :-)

OpenStudy (bibby):

that's not helping. you fed her the answer

OpenStudy (anonymous):

I didn't really understand the way you put it. She/he might have given me the answer, but I understood the way she got it.